OpenStudy (gabylovesyou):

Reggie can line a football field in 120 minutes. Rosalinda can line a football field in 80 minutes. If they work together, how many minutes does it take them to line a football field? 40 minutes 48 minutes 200 minutes 240 minutes

OpenStudy (gabylovesyou):

@jim_thompson5910

OpenStudy (charles.305):

I'd say it is probably be 200 minutes.

jimthompson5910 (jim_thompson5910):

It's very similar to the last problem x = time it takes Reggie to do the job alone (with no help) y = time it takes Rosalinda to do the job alone (with no help) z = time it takes to do the job if the two worked together All times (x,y,z) are in minutes In this case, x = 120 min y = 80 min z = unknown \[\Large \frac{1}{x} + \frac{1}{y} = \frac{1}{z}\] \[\Large \frac{1}{120} + \frac{1}{80} = \frac{1}{z}\] solve for z

OpenStudy (aihberkhan):

Okay. This link should help you. http://brainly.com/question/636217

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @Gabylovesyou